33. QC In a test run, a certain car accelerates uniformly Qlc from zero to 24.0 m/s in 2.95 s. (a) What is the mag- nitude of the cars acceleration? (b) How long does it take the car to change its speed from 10.0 m/s to 20.0 m/s? (c) Will doubling the time always double the change in speed? Why?
Answer #1

(A) vf = vi + a t

24 = 0 + 2.95a

a = 8.14 m/s^2 .....Ans

(B) 20 = 10 + 8.14t

t = 1.23 sec

(C) change in sped = vf - vi = a t

a is constant hence doubling the time will double the change in speed.